OpenStudy (anonymous):

the unit of acceleration is "m/s2". why it is "s2"????

Parth (parthkohli):

Hi. Acceleration is about how much the velocity increases in some time. \[\rm acceleration = \dfrac{change~in ~velocity}{time}\]The change in velocity has the same units as that of velocity, here is \(m/s\). Time has the units of time, which obviously here would be \(s\). So acceleration has the units\[\dfrac{m/s}{s} = \dfrac{m}{s \cdot s} = \dfrac{m}{s^2}\]

OpenStudy (anonymous):